The following data give the prices of seven textbooks randomly selected from a university bookstore. $92 $176 $102 $126 $55 $155 $148 a. Find the mean for these data. Calculate the deviations of the data values from the mean. Is the sum of these deviations zero? Mean = $ 122 Deviation from the mean for $176- sum of these deviations b. Calculate the range, variance, and standard deviation. [Round your answers to 2 decimal places.] Range = $ Variance Standard deviation $
Answer #1

(a)

(i)

Mean (ar{x}) is given by:

854 = 122 41=854

(ii)

Deviation from the mean for $176 = 176 - 122 = 54

(iii)

To find the sum of these deviations:

x x - ar{x}
92 - 30
176 54
102 - 20
126 4
55 - 67
155 33
148 26
Total = 0

Thus, we get
Sum of these deviations = 0

(b)

(i)

Range = Maximum -Minimum = 176 - 55 = 121.00

(ii)

To find Variance:

x x - ar{x} (x - ar{x})2
92 - 30 900
176 54 2916
102 - 20 400
126 4 16
55 - 67 4489
155 33 1089
148 26 676
Total = 0 10486

Variance (s2) is given by:

Σί= 1 (zi-i)2 6 10486 6 1747.67

(iii)

Standard deviation (s) is given by:

s- V1747.67-41.81